OpenStudy (meilendurcer):

-5/3 divide -15/7= A. 7/9 B. -3 4/7 C. -7/9 D. 2 2/15

OpenStudy (anonymous):

division is multiplication by reciprocal so : (-5/3) / (-15/7) = (-5/3) * (7/-15) can you do it now ?

OpenStudy (anonymous):

for this multiply the first rational by the multiplicative inverse of the second, cancel out common factors and see what you get

OpenStudy (meilendurcer):

@Coolsector 3.424?

OpenStudy (anonymous):

no.. lets do this together : (-5/3) / (-15/7) = (-5/3) * (7/-15) = (-35 / -45) = 7/9

OpenStudy (anonymous):

did you use a calculator?????

OpenStudy (meilendurcer):

yes... and Thank you

OpenStudy (anonymous):

-5/3 divide -15/7= -5 7 -35 7 ---- * ----- = ------- = ----- 3 -15 -45 9